-13m-(-4n)-12=19
n-6m+(-13)=-8
M=
N=

Answer by NirMath4U(7)

-13m+4n=19+12
n=6m+13-8
n=6m+5
-13m+4(6m+5)=31
-13m+24m+20=31
11m=31-20
11m=11
m=1
n=6*1+5=11